Your friend is asking you to invest in his bakery, and you have 35000. The returns are volatile, and you may get either 43000 with a probability of 0.44 or 27000 with a probability of 0.56.



a. What is the expected value of the investment?



b. Will you buy the refrigerator? Why?



c. Identify the mean, variance, and standard deviation.
1
Expert's answer
2021-10-04T18:59:16-0400

Solution:

a.

The expected value of the investment"=43000 \times 0.44 + 27000 \times 0.56"

"=18920+15120
\\=34040"

b. I have 35000. I may get an expected value of 34040, which is less than the amount I have to invest. I will not buy the refrigerator as it provides me the loss.

c.

Mean"=E[X]=34040"

"E[X^2]=43000^2\times0.44+27000^2\times0.56=1221800000"

Now, variance"=\sigma^2=E[X^2]-(E[X])^2"

"=1221800000-34040^2
\\=63078400"

Standard deviation"=\sigma=7942.191"